suppose c=150+0.8(Y-t), T=100+0.1Y, G=300 and I= 210-100r derive the equation of IS curve and find out its slope. Suppose G changes by 5 units how much will the IS shift.
1
Expert's answer
2020-11-26T07:22:54-0500

"Y=C+I+G\newline =150+0.8(Y-t)+210-100r+300\newline =150+0.8Y-0.8t+210+100r+300\newline Y= 660+0.8Y -0.8t+100\newline Y-0.8Y=660-0.8t +100r\newline 0.2Y= 660-0.8t+100r\newline Y=3300-4t +500r"